OpenStudy (anonymous):

Prove that the following equation is an identity: (six^3y-cos^3y)/siny-cosy = (2+sin2y)/2

OpenStudy (mertsj):

\[\frac{\sin ^3y-\cos ^3y}{\sin y-\cos y}=\frac{(\sin y-\cos y)(\sin ^2y+\sin y \cos y+\cos ^2y)}{\sin y-\cos y}\]= \[\frac{(\sin y-\cos y)(1+\sin y \cos y)}{\sin y-\cos y}=(1+\sin y \cos y)\times\frac{2}{2}=\] \[\frac{2+2\sin y \cos y}{2}=\frac{2+\sin 2y}{2}\]

OpenStudy (anonymous):

Thanks.

OpenStudy (mertsj):

yw